Any permanent decrease in autonomous real spending will cause even larger decreases in real GDP per year.

A. False

B. True

If the MPC​ = 0.8, a permanent increase in planned real investment of​ $40 billion will increase real GDP by a total of

A. $200 billion

B. $40 billion.

C. $80 billion.

D. $160 billion.
Answer #1

1. Ans: True

Explanation:

when autonomous expenditure decreases, real GDP will decrease by a larger amount due to multiplier effect. Thus, the statement is true.

2. Ans: $200 billion

Explanation:

Multiplier = 1 / (1 - MPC) = 1 / (1 - 0.8) = 5

Increase in total real GDP = 5 * $40 billion = $200 billion